# Linear Bandits with Partially Observable Features

### v2

@@ G.5 A Bound for the Probability of Exploration
For each $t$, let $\widehat{a}_{t}=\mathop{\mathrm{argmax}}_{a\in[K]}\widetilde{\mathbf{x}}_{a}^{\top}\widehat{\boldsymbol{\mu}}_{t}$ denote the maximizing action based on the estimator $\widehat{\boldsymbol{\mu}}_{t}$. Then the action $a_{t}$ chosen by the resampling scheme in Algorithm 1 and Algorithm 2 satisfies,
@@
Since the algorithm resampling at most $\rho_{t}$ times, for a fixed $t\in[T]$,
$\mathbb{P}(a_{t}\neq\hat{a}_{t})$ =($=\mathbb{P}\left(\{\widetilde{a}_{t}=a_{t}\}\cap\{a_{t}\neq\hat{a}_{t}\}\right)+\mathbb{P}\left(\{\widetilde{a}_{t}\neq a_{t}\}\cap\{a_{t}\neq\hat{a}_{t}\}\right)$
($\leq\mathbb{P}\left(\{\widetilde{a}_{t}=a_{t}\}\cap\{a_{t}\neq\hat{a}_{t}\}\right)+\mathbb{P}(\widetilde{a}_{t}\neq a_{t})$
=($=\mathbb{P}\left(\{\widetilde{a}_{t}=a_{t}\}\cap\{a_{t}=k\}\right)+\underbrace{\mathbb{P}(\widetilde{a}_{t}\neq a_{t})}_{\text{Failure of resampling}}\quad\text{for }k\neq\hat{a}_{t},$ (27)
where $\mathbb{P}(a_{t}=k)=t^{-1/2}/(K-1)$ and $\mathbb{P}(\widetilde{a}_{t}=a_{t})=p$, defined by Algorithm 1 and Eq. 9, respectively. For the first term in Eq. 27, by union bound it holds that:
($\mathbb{P}\left(\{\widetilde{a}_{t}=a_{t}\}\cap\{a_{t}=k\}\right)$ $=\mathbb{P}\left(\bigcup_{m=1}^{\rho_{t}}\{\text{Resampling success at trial }m\}\cap\{a_{t}=k\}\right)$
$\leq\sum_{m=1}^{\rho_{t}}\mathbb{P}\left(\{\text{Resampling success at trial }m\}\cap\{a_{t}=k\}\right),$
@@
$=\frac{1}{\sqrt{t}(K-1)}.$
which, combined with Eq. 27, results in:

### v3

@@ G.5 A Bound for the Probability of Exploration
For each $t$, let $\widehat{a}_{t}=\mathop{\mathrm{argmax}}_{a\in[K]}\widetilde{\mathbf{x}}_{a}^{\top}\widehat{\boldsymbol{\mu}}_{t}$ denote the maximizing action based on the estimator $\widehat{\boldsymbol{\mu}}_{t}$. Then the action $a_{t}$ chosen by the resampling scheme in Algorithm˜1 and Algorithm˜2 satisfies,
@@
Since the algorithm resamples at most $\rho_{t}$ times, for a fixed $t\in[T]$,
$\mathbb{P}(a_{t}\neq\hat{a}_{t})$ $=\mathbb{P}\left(\{\widetilde{a}_{t}=a_{t}\}\cap\{a_{t}\neq\hat{a}_{t}\}\right)+\mathbb{P}\left(\{\widetilde{a}_{t}\neq a_{t}\}\cap\{a_{t}\neq\hat{a}_{t}\}\right)$
$\leq\mathbb{P}\left(\{\widetilde{a}_{t}=a_{t}\}\cap\{a_{t}\neq\hat{a}_{t}\}\right)+\mathbb{P}(\widetilde{a}_{t}\neq a_{t})$
$=\mathbb{P}\left(\{\widetilde{a}_{t}=a_{t}\}\cap\{a_{t}=k\}\right)+\underbrace{\mathbb{P}(\widetilde{a}_{t}\neq a_{t})}_{\text{Failure of resampling}}\quad\text{for }k\neq\hat{a}_{t},$ (36)
where $\mathbb{P}(a_{t}=k)=t^{-1/2}/(K-1)$ and $\mathbb{P}(\widetilde{a}_{t}=a_{t})=p$, defined by Algorithm˜1 and Eq.˜9, respectively. For the first term in Eq.˜36, by applying the union bound, we obtain
$\mathbb{P}\left(\{\widetilde{a}_{t}=a_{t}\}\cap\{a_{t}=k\}\right)$ $=\mathbb{P}\left(\bigcup_{m=1}^{\rho_{t}}\{\text{Resampling success at trial }m\}\cap\{a_{t}=k\}\right)$
$\leq\sum_{m=1}^{\rho_{t}}\mathbb{P}\left(\{\text{Resampling success at trial }m\}\cap\{a_{t}=k\}\right)$
@@
$=\frac{1}{\sqrt{t}(K-1)},$
which, combined with Eq.˜36, gives